AN EFFICIENT ARCHITECTURE FOR SOLVING THE RECURSIVE CONVOLUTION EQUATION WITH HIGH THROUGHPUT

B. B. ZHOU AND RICHARD P. BRENT

Abstract

The one-dimensional (1D) recursive convolution problem plays a very important role in digital signal processing. It is usually considered that recursion implies sequential execution. Thus only 1D architectures were taken into account in VLSI implementation of this kind of problem. In this paper, we derive a two-dimensional systolic architecture for solving this problem with a throughput rate of two. Our solution has about twice the area, but twice the throughput of the previously known 1D systolic arrays. We also give an example to show that, by using our systolic architecture, IIR problems can be computed systolically with a throughput rate of two.

Comments

Only the Abstract is given here. The full paper appeared as [1]. For related work, see [2, 3, 4].

References

1. B. B. Zhou and R. P. Brent, “An efficient architecture for solving the recursive convolution equation with high throughput”, Proc. First IASTED International Symposium on Signal Processing and its Applications (edited by B. Boashash), Institution of Engineers, Australia, Vol. 2, 1987, 771–775. ISBN 0-8581-4143-4. rpb099.

2. B. B. Zhou and R. P. Brent, “A high throughput systolic implementation of the second order recursive filter”, Proceedings IEEE 1988 Conference on Acoustics, Speech and Signal Processing (New York, April 1988). Also appeared as Report TR-CS-88-04, Computer Sciences Laboratory, ANU, February 1988, 14 pp. rpb103.

3. B. B. Zhou and R. P. Brent, A two-level pipelined implementation of direct-form recursive filters, Report TR-CS-88-06, Computer Sciences Laboratory, ANU, April 1988, 16 pp. rpb104.

4. R. P. Brent and B. B. Zhou, “A stabilized parallel algorithm for direct-form recursive filters”, IEEE Transactions on Computers 40 (1991), 333–336. Also appeared as A stabilized parallel implementation of direct-form recursive filters, Report TR-CS-88-07, Computer Sciences Laboratory, ANU, May 1988, 9 pp. rpb105.

Computer Sciences Laboratory, Australian National University, Canberra, ACT 0200 E-mail address: {bing,rpb}@cslab.anu.edu.au

1991 Mathematics Subject Classification. Primary 65Q05; Secondary 65G05. Key words and phrases. Filter, IIR problem, recursive filter, recursive convolution equation, direct-

form filter, stability, parallel algorithm, pipeline, systolic architecture, VLSI.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1987, IEAust. Comments \begin{matrix}c \\ ◯ \end{matrix} 1993, R. P. Brent.

rpb099a typeset using AMS-L^{A}TEX